# An Extensible Software Transport Layer for GPU Networking

### v1

@@ Abstract.
Fast-evolving machine learning (ML) workloads have increasing requirements for networking. However, host network transport on RDMA NICs is hard to evolve, causing problems for ML workloads. For example, single-path RDMA traffic is prone to flow collisions that severely degrade collective communication performance. We present UCCL, an extensible software transport layer to evolve GPU networking. UCCL decouples the data path and control path of existing RDMA NICs and efficiently runs the control-path transport on host CPUs. This software extensibility brings in transport innovations that cannot be achieved in hardware for ML workloads, e.g., a multipath transport to resolve flow collisions. ML collectives atop UCCL achieve up to 3.3 higher performance compared to an industry solution.
@@ 1. Introduction
To address the second challenge, UCCL leverages an array of techniques tailored to the unique characteristics of ML workloads. On the one hand, UCCL leverages GPUDirect (Corporation, 2025d) to alleviate CPU overhead. On the other hand, UCCL employs control coalescing to make transport control decisions for every 32KB data chunk rather than every packet; this works well as many transport decisions such as CC do not require per-packet reaction, but just per-RTT to avoid overreaction (Li et al., 2019; Google, 2024). As a result, UCCL can handle 400 Gbps unidirectional traffic with a single CPU core, and achieve the same message latency as hardware-based transport. UCCL leverages 256 RDMA QPs (Queue Pairs) per connection to perform multipathing, without worrying about the high QP context swapping overhead highlighted in previous work (Wang et al., 2023); this is because ML workloads feature bulk data transfer with mostly MTU-sized large packets, which effectively amortizes the swapping overhead. UCCL’s software transport is practical and economical in modern GPU servers, which usually have hundreds of powerful CPU cores (in fact, these CPU cores are often heavily underutilized).
UCCL provides an expressive yet easy-to-use interface. To demonstrate the versatility of this interface and the power of UCCL’s extensibility, we use three case studies. First, we implement a multipath transport protocol that mitigates flow collisions by leveraging packet spraying, i.e., randomly sending packets from a single connection across different paths (Dixit et al., 2013). This transport implementation achieves 3.3 higher throughput for collective communication over AWS’ existing solution, SRD (Shalev et al., 2020) on EFA NICs, which translates into up to 1.4 speedup for two ML applications. Second, we implement a receiver-driven protocol EQDS (Olteanu et al., 2022) to handle network incast in MoE-like workloads, achieving 4.9 better message tail latency over InfiniBand built-in transport. Third, we implement selective retransmission (Mathis et al., 1996) for efficient transport loss recovery, and show its superiority over RDMA hardware transport under packet loss. These case studies highlight that UCCL can effectively enable innovations in the transport protocols that are hard to implement in today’s network stack without expensive and time-consuming changes.
@@ 2.1. GPU Networking: from RDMA to Collectives
GPU networking is extremely heterogeneous. At the high level, GPU collective communication libraries like NVIDIA NCCL (NVIDIA Corporation, 2025) and AMD RCCL (AMD, 2025) use RDMA and kernel TCP (non-RDMA) for inter-server networking, with RDMA preferred since it is faster and more efficient. RDMA provides various communication primitives called Queue Pairs (QPs) including Reliable Connection (RC), Unreliable Connection (UC), and Unreliable Datagram (UD):
@@
Figure 1 shows an overview of how collective communication uses RDMA. Once the ML application calls a collective like allreduce, the collective library launches a reduction kernel on each participant GPU to handle data reduction/copy. Next, the sender CPU issues an RDMA write over RC QPs to transfer data. The receiver CPU polls a completion flag in its memory, which the sender sets upon write completion. The library manages a set of transport buffers on the GPU memory to buffer the RDMA data, and relies on the GPU kernel to copy data between the transport buffers and application tensor buffers. The GPU kernel also performs reductions (e.g., sum, max) on the transport buffers (from multiple sender GPUs) into the tensor buffers.
@@ 2.2. Motivation on Extensibility
Receiver-driven CC for incast. Recent MoE serving workloads are prone to network incast problems. In DeepSeek’s online deployment of their V3 model (Liu et al., 2024), each of the 320 GPUs holds a single expert module, where hidden states are exchanged between expert modules across GPUs, i.e., Expert Parallelism (EP). As the request pattern and load change over time, some experts become much hotter than others, receiving more network traffic from other experts leading to network incast issues. DeepSeek reports the hottest expert could receive 10 more load than the average one. This network incast can be better handled by receiver-driven CC (Handley et al., 2017; Olteanu et al., 2022) which controls last-hop congestion—unfortunately, there is no receiver-driven CC on commercial RDMA NICs.
@@
Heterogeneous NICs. Datacenters usually consist of multiple generations and vendors of RDMA NICs due to continuous expansion, cost optimization, and to avoid vendor lock-in. While NVIDIA, Broadcom, and AMD all have 400 Gbps RDMA NICs for ML (NVIDIA Corporation, 2024a; Broadcom, 2024; AMD Pensando, 2024), they come with subtly different control path logic like packet reliability and CC; this heterogeneity reduces achievable bandwidth by 2-33 when communicating between NICs from different generations/vendors, as reported by Alibaba (Li et al., 2023). Instead, if we can extensibly align these NICs’ control path logic in software, we could avoid such a severe performance drop.
@@ 2.3. Prior Work on Extensibility
Leveraging SmartNICs. Several recent efforts aim to make RDMA transport programmable by offloading it to the SmartNIC RISC cores, but they have constraints in extensibility and performance. Google Falcon SmartNICs (Google, 2024, 2024) only support programming rate update actions for latency-based Swift CC (Kumar et al., 2020) and path selection decisions with limited paths (Qureshi et al., 2022); similar constraints also apply to firmware updates on hardware RDMA NICs. AMD Pensando SmartNICs (AMD Pensando, 2024) support using P4 language (Bosshart et al., 2014) to program their transport layer, but P4 has limited programmability, e.g., hard to implement efficient loss recovery; it is also unclear if they could support receiver-driven CC. AWS EFA SmartNICs (Services, 2025) implement a proprietary multipath reliable transport SRD (Shalev et al., 2020) using NIC ARM cores. However, they perform poorly on collectives, especially for the connection-intensive all-to-all (see §6.1). We attribute this to the fundamentally limited processing power and cache capacity of these wimpy RISC cores (Phothilimthana et al., 2018; Liu et al., 2019). FPGA-based SmartNICs provide higher performance but have limited extensibility due to hardware resource constraints (Wang et al., 2023).
@@
How to separate control and data (§3.1)? | How to reassemble packets on receivers (§3.1)? | How to enable multipathing (§3.2)? | How to build efficient software transport (§3.3)? | |
RC, UC | Write with immediate | RDMA NIC handles | Different QP connections | Control coalescing |
UD | Send/Recv with scatter-gather | GPU does scattered memcpy | Src-dst QP combinations | Chained posting |
AF_XDP | – | CPU handles | Different UDP ports | User-space packet IO |
@@ 3. UCCL Design
In UCCL, all connections between a specific pair of NICs share the same set of QPs (e.g., 256), including the case of multiple GPUs sharing one NIC. This design fully harnesses the underlying multipath datacenter network while not burning too many QPs (§3.2). UCCL further integrates a bag of techniques to run software transport as efficiently as possible, e.g., control coalescing, connection splitting, chained posting, and more (§3.3). UCCL also supports extending transport for non-RDMA NICs through the AF_XDP user-space packet IO, bypassing conventional kernel TCP stacks. Table 1 summarizes the UCCL techniques to achieve transport extensibility across different communication primitives. We now describe them in the rest of this section.
@@ 3.1. Separating Control Path and Data Path
UC as the preferred QP. UCCL chooses UC as the preferred QP whenever available on the RDMA NIC, similar to Flor (Li et al., 2023), because it supports efficient segmentation and reassembly offloading to the NIC (in contrast to UD), while bypassing hardware-baked CC, loss recovery, and out-of-order packet handling (in contrast to RC). Under UC, UCCL uses the efficient RDMA verb of write with immediate to transfer data chunks; this verb operates in a two-sided mode, so that both sender and receiver CPUs can react to the data transfers. For the data path, the sender CPU specifies the addresses of source and destination data chunks when issuing the verb, and then the sender NIC will automatically segment the source data chunk into MTU-sized packets with packet headers prepended and send them out. When receiving these packets, the receiver NIC will remove the packet headers and reassemble payloads into the contiguous memory region specified by the sender CPU.
For the control path, write with immediate allows carrying a 32-bit imm_data from the sender CPU to the receiver CPU, serving as the control header for UCCL transport. UC guarantees that any successfully arrived chunk will generate a CQE with the imm_data embedded, which is then consumed by the receiver CPU. Within the 32-bit budget, UCCL allocates 8 bits for connection ID and 7 bits for message ID, supporting 256 connections for each pair of NICs on a UCCL engine and 128 in-flight messages for each connection, which is sufficient for collective communication. UCCL then allocates 8 bits to the chunk sequence number (CSN) to identify the position of a chunk in the message being transferred. Another 1 bit is allocated for marking the last chunk of a message. The remaining 8 bits are reserved for more advanced CC such as receiver-driven ones (see §4.2).
@@ 3.2. Harnessing Multipath
One of the key motivations for GPU network extensibility is to harness the multipath capacity of modern datacenter networks (§2.2). UCCL achieves this by using multiple UC, RC, or UD QPs. Basically, network traffic from different QPs will likely go through different network paths, as both RoCE and Infiniband usually use ECMP (Equal-Cost Multi-Path) for multipath routing with source and destination QP numbers as the hash inputs (Gangidi et al., 2024; NVIDIA Corporation, 2024b). For UC and RC, UCCL by default uses 256 QPs, which provides maximum 256 different network paths as used by recent transport research (Le et al., 2024; Bonato et al., 2024). For UD, UCCL uses a much smaller number of QPs by combining different source and destination QPs. For example, 16 source UD QPs and 16 destination UD QPs will provide maximum 1616=256 different network paths, because for connection-less UD, each source QP can send packets to any destination QP. UCCL also supports a configurable number of QPs for different collectives, e.g., a smaller number might work well for all-to-all with relatively high entropy (Gangidi et al., 2024). To avoid burning too many QPs, especially when the collective library creates multiple connections between the same pair of NICs (e.g., because of multiple GPUs sharing one NIC), UCCL lets all these connections share the same set of QPs.
We note that making this multi-QP design choice is not trivial, especially as a line of prior work has highlighted the severe QP scalability issues on RDMA NICs (Kalia et al., 2016, 2019; Chen et al., 2019; Monga et al., 2021; Kong et al., 2022; Wang et al., 2023; Lee et al., 2024). For example, SRNIC (Wang et al., 2023) reports 23% bandwidth drop when scaling RC QPs from 256 to 512, and 46% drop when scaling to 16k. This bandwidth drop is caused by the QP swapping overhead: the NIC can only hold/cache limited QP contexts on its SRAM, and must spill/swap the excessive rest to the host DRAM over PCIe, incurring frequent QP swapping. Surprisingly, we do not observe such a severe performance drop for collective communication—we only see 17% drop when scaling RC QPs to 60k (see §6.4.1); for UC with a smaller QP context size, the drop is even negligible.
@@ 3.3. Towards Efficient Software Transport
So far, we have discussed how UCCL decouples the control path and data path to make flexible transport decisions on the CPU, and how UCCL achieves multipathing. The next question is how to efficiently implement a software multipath transport to support the high bandwidth in GPU networking. This is challenging as a single GPU server could have 8400 Gbps RDMA NICs, totaling 3.2 Tbps bandwidth bidirectionally (Huang et al., 2012); the next generation RDMA NIC will achieve 800 Gbps (Corporation, 2025c), rendering 6.4 Tbps bandwidth. As a reference, Google’s software transport Snap (Marty et al., 2019) can handle 80 Gbps traffic on a CPU core (though they do not use RDMA NICs). Our goal is to use 1 CPU core to handle 400G unidirectional traffic. To this end, we leverage the following techniques:
@@
Small message bypassing multipath. Using 256 UC/RC QPs for small messages still incurs a high performance penalty in terms of latency; this is because these small message transfers cannot fully amortize the QP swapping overhead. To optimize this, UCCL empirically bypasses multipathing for messages under 8KB and transmits them directly via a dedicated UC/RC QP. We argue this bypassing design for small messages does not impact the overall benefits of multipathing, which usually begins to matter at high traffic load with large message transfers (Le et al., 2024).

### v2

@@ Abstract.
Fast-evolving machine learning (ML) workloads have increasing requirements for networking. However, host network transport on RDMA NICs is hard to evolve, causing problems for ML workloads. For example, single-path RDMA traffic is prone to flow collisions that severely degrade collective communication performance. We present UCCL, an extensible software transport layer to evolve GPU networking. UCCL decouples the data path and control path of existing RDMA NICs and efficiently runs the control-path transport on host CPUs. This software extensibility brings in transport innovations that cannot be achieved in hardware for ML workloads, e.g., a multipath transport to resolve flow collisions. ML collectives atop UCCL achieve up to 4.5 higher performance compared to existing RDMA NICs.
@@ 1. Introduction
To address the second challenge, UCCL leverages an array of techniques tailored to the unique characteristics of ML workloads. On the one hand, UCCL leverages GPUDirect (Corporation, 2025f) to alleviate CPU overhead. On the other hand, UCCL employs control coalescing to make transport control decisions for every 32KB data chunk rather than every packet; this works well as many transport decisions such as CC do not require per-packet reaction, but just per-RTT to avoid overreaction (Li et al., 2019; Google, 2024). As a result, UCCL can handle 400 Gbps unidirectional traffic with a single CPU core, and achieve the same message latency as hardware-based transport. UCCL leverages 256 RDMA QPs (Queue Pairs) per connection to perform multipathing, without worrying about the high QP context swapping overhead highlighted in previous work (Wang et al., 2023); this is because ML workloads feature bulk data transfer with mostly MTU-sized large packets, which effectively amortizes the swapping overhead. UCCL’s software transport is practical and economical in modern GPU servers, which usually have hundreds of powerful CPU cores; in fact, these CPU cores are often heavily underutilized. For example, from our private conversation with the model training team of a major GPU vendor, their CPU utilization is averagely 14.5% out of 128 cores, when using an internal version of Megatron-LM (NVIDIA Corporation, 2025a) for model training.
UCCL provides an expressive yet easy-to-use interface. To demonstrate the versatility of this interface and the power of UCCL’s extensibility, we use three case studies. First, we implement a multipath transport protocol that mitigates flow collisions by leveraging packet spraying, i.e., randomly sending packets from a single connection across different paths (Dixit et al., 2013). Compared to existing RDMA transport on hardware, ML collectives using our transport achieve up to 4.5 higher throughput on NVIDIA ConnectX-7 NICs, and up to 1.9 higher on Broadcom Thor-2 NICs with rail-optimized topology. Second, we implement a receiver-driven protocol EQDS (Olteanu et al., 2022) to handle network incast in MoE-like workloads, achieving 4.9 better message tail latency over InfiniBand built-in transport. Third, we implement selective retransmission (Mathis et al., 1996) for efficient transport loss recovery, and show its superiority over RDMA hardware transport under packet loss. These case studies highlight that UCCL can effectively enable innovations in the transport protocols that are hard to implement in today’s network stack without expensive and time-consuming changes. UCCL is open-sourced at https://github.com/uccl-project/uccl.
@@ 2.1. GPU Networking: from RDMA to Collectives
GPU networking is extremely heterogeneous. At the high level, GPU collective communication libraries like NVIDIA NCCL (NVIDIA Corporation, 2025b) and AMD RCCL (AMD, 2025) use RDMA and kernel TCP (non-RDMA) for inter-server networking, with RDMA preferred since it is faster and more efficient. RDMA provides various communication primitives called Queue Pairs (QPs) including Reliable Connection (RC), Unreliable Connection (UC), and Unreliable Datagram (UD):
@@
Figure 1 shows an overview of how collective communication uses RDMA. Once the ML application calls a collective like allreduce, the collective library launches a reduction kernel on each participant GPU to handle data reduction/copy. Next, the sender CPU issues multiple RDMA writes over RC QPs to transfer data chunk by chunk. The receiver CPU polls completion flags in its memory, which the sender sets upon write completions. The library manages a set of transport buffers on the GPU memory to buffer the RDMA data, and relies on the GPU kernel to copy data between the transport buffers and application tensor buffers. The GPU kernel also performs reductions (e.g., sum, max) on the transport buffers (from multiple sender GPUs) into the tensor buffers.
@@ 2.2. Motivation on Extensibility
Receiver-driven CC for incast. Recent MoE serving workloads are prone to network incast problems. In DeepSeek’s online deployment of their 671B V3 model (Liu et al., 2024), each of the 320 GPUs holds a single expert module, where hidden states are exchanged between expert modules across GPUs, i.e., Expert Parallelism (EP). As the request pattern and load change over time, some experts become much hotter than others, receiving more network traffic from other experts leading to network incast issues. DeepSeek reports the hottest expert could receive 10 more load than the average one. Expert load balancing algorithms such as EPLB (DeepSeek AI, 2025) try to balance the load by dynamically replicating experts. However, this happens at a much slower pace (e.g., 10 minutes (DeepSeek, 2025)) to avoid the high cost of moving experts, thus unable to handle transient incast. Such transient network incast can be better handled by receiver-driven CC (Handley et al., 2017; Olteanu et al., 2022), which controls last-hop congestion—unfortunately, there is no receiver-driven CC on commercial RDMA NICs.
@@
Heterogeneous NICs. Datacenters usually consist of multiple generations and vendors of RDMA NICs due to continuous expansion, cost optimization, and to avoid vendor lock-in. While NVIDIA, Broadcom, AMD, and more vendors all have 400 Gbps RDMA NICs for ML (NVIDIA Corporation, 2024a; Broadcom, 2024; AMD Pensando, 2024), they come with subtly different control path logic such as packet reliability and CC. In practice, this heterogeneity reduces achievable bandwidth by 2-33 when communicating between NICs from different generations/vendors, as reported by Alibaba (Li et al., 2023). Prior work Flor (Li et al., 2023) has shown that extensibly aligning these NICs’ control path logic in software could avoid such a severe performance drop.
@@ 2.3. Prior Work on Extensibility
Leveraging SmartNICs. Several recent efforts aim to make RDMA transport programmable by offloading it to the SmartNIC RISC cores, but they have constraints in extensibility and performance. Google Falcon SmartNICs (Google, 2024, 2024) only support programming rate update actions for latency-based Swift CC (Kumar et al., 2020) and path selection decisions with limited paths (Qureshi et al., 2022); similar constraints also apply to firmware updates on hardware RDMA NICs. AMD Pensando SmartNICs (AMD Pensando, 2024) support using P4 language (Bosshart et al., 2014) to program their transport layer, but P4 has limited programmability, e.g., hard to implement efficient loss recovery; it is also unclear if they could support receiver-driven CC. FPGA-based SmartNICs provide higher performance but have limited extensibility due to hardware resource constraints (Wang et al., 2023). AWS EFA SmartNICs (Services, 2025) implement a proprietary multipath reliable transport SRD (Shalev et al., 2020) with out-of-order packet delivery using NIC ARM cores, to address network congestion in HPC and ML workloads. The SRD protocol is implemented using EFA-specific firmware and supports live upgrade with good extensibility. However, we empirically find that the EFA SmartNICs on AWS p4d.24xlarge GPU VMs perform poorly for connection-intensive all-to-all collectives (see §6.1). We attribute this to the limited processing power and cache capacity of SmartNIC ARM cores due to power constraints, as also demonstrated by prior research on publicly available SmartNICs (Phothilimthana et al., 2018; Liu et al., 2019; Schuh et al., 2021).
Note that the above all-to-all measurement uses the AWS EFA NICs on the p4d.24xlarge instances, and thus it may not apply to their new-generation EFA NICs on p5/p5en/p6 instances. Specifically, the ever-changing EFA firmware and upgraded EFA hardware may lead to different performance results for connection-intensive all-to-all collectives. In general, if these SmartNICs get upgraded with higher processing power and cache capacity to better handle all-to-all, we view it as an echo to UCCL’s high-level methodology of software extensibility for GPU networking.
@@ 3. UCCL Design
In UCCL, all connections between a specific pair of NICs share the same set of QPs (e.g., 256), including the case of multiple GPUs sharing one NIC. This design fully harnesses the underlying multipath datacenter network while not burning too many QPs (§3.2). UCCL further integrates a bag of techniques to run software transport as efficiently as possible, e.g., control coalescing, connection splitting, chained posting, and more (§3.3). UCCL also supports extending transport for non-RDMA NICs through the AF_XDP user-space packet IO, bypassing conventional kernel TCP stacks. We now describe them in the rest of this section.
@@ 3.1. Separating Control Path and Data Path
UC as the preferred QP. UCCL chooses UC as the preferred QP whenever available on the RDMA NIC, similar to Flor (Li et al., 2023), because it supports efficient segmentation and reassembly offloading to the NIC (in contrast to UD), while bypassing hardware-baked CC, loss recovery, and out-of-order packet handling (in contrast to RC). As shown in Figure 3, UCCL uses the efficient RDMA verb of write with immediate to transfer data chunks over UC; this verb operates in a two-sided mode, so that both sender and receiver CPUs can react to the data transfers. For the data path, the sender CPU specifies the addresses of source and destination data chunks when issuing the verb, and then the sender NIC will automatically segment the source data chunk into MTU-sized packets with packet headers prepended and send them out. When receiving these packets, the receiver NIC will remove the packet headers and reassemble payloads into the contiguous memory region specified by the sender CPU.
For the control path, write with immediate allows carrying a 32-bit imm_data from the sender CPU to the receiver CPU, serving as the control header for UCCL transport. Figure 3 shows an example. UC guarantees that any successfully arrived chunk will generate a CQE with the imm_data embedded, which is then consumed by the receiver CPU. Within the 32-bit budget, UCCL allocates 8 bits for connection ID and 7 bits for message ID, supporting 256 connections for each pair of NICs on a UCCL engine and 128 in-flight messages for each connection, which is sufficient for collective communication. UCCL then allocates 8 bits to the chunk sequence number (CSN) to identify the position of a chunk in the message being transferred. Another 1 bit is allocated for marking the last chunk of a message. The remaining 8 bits are reserved for more advanced CC such as receiver-driven ones (see §4.2).
@@ 3.2. Harnessing Multipath
One of the key motivations for GPU network extensibility is to harness the multipath capacity of modern datacenter networks (§2.2). UCCL achieves this by using multiple UC, RC, or UD QPs, as shown in Figure 5. Basically, network traffic from different QPs will likely go through different network paths, as both RoCE and Infiniband usually use ECMP (Equal-Cost Multi-Path) for multipath routing with source and destination QP numbers as the hash inputs (Gangidi et al., 2024; NVIDIA Corporation, 2024c). For UC and RC, UCCL by default uses 256 QPs, which provides maximum 256 different network paths as used by recent transport research (Le et al., 2024; Bonato et al., 2024). For UD, UCCL uses a much smaller number of QPs by combining different source and destination QPs. For example, 16 source UD QPs and 16 destination UD QPs will provide maximum 1616=256 different network paths, because for connection-less UD, each source QP can send packets to any destination QP. UCCL also supports a configurable number of QPs for different collectives, e.g., a smaller number might work well for all-to-all with relatively high entropy (Gangidi et al., 2024). To avoid burning too many QPs, especially when the collective library creates multiple connections between the same pair of NICs (e.g., because of multiple GPUs sharing one NIC), UCCL lets all these connections share the same set of QPs.
We note that making this multi-QP design choice is not trivial, especially as a line of prior work has highlighted the severe QP scalability issues on RDMA NICs (Kalia et al., 2016, 2019; Chen et al., 2019; Monga et al., 2021; Kong et al., 2022; Wang et al., 2023; Lee et al., 2024). For example, SRNIC (Wang et al., 2023) reports 23% bandwidth drop when scaling RC QPs from 256 to 512, and 46% drop when scaling to 16k. This bandwidth drop is caused by the QP swapping overhead: the NIC can only hold/cache limited QP contexts on its SRAM, and must spill/swap the excessive rest to the host DRAM over PCIe, incurring frequent QP swapping. Surprisingly, we do not observe such a severe performance drop for collective communication—Figure 6 shows only 17% drop when scaling RC QPs from 60 to 60k, and a negligible drop for UC (that has a smaller QP context size than RC).
@@ 3.3. Towards Efficient Software Transport
So far, we have discussed how UCCL decouples the control path and data path to make flexible transport decisions on the CPU, and how UCCL achieves multipathing. The next question is how to efficiently implement a software multipath transport to support the high bandwidth in GPU networking. This is challenging as a single GPU server could have 8400 Gbps RDMA NICs, totaling 3.2 Tbps bandwidth bidirectionally (Huang et al., 2012); the next generation RDMA NIC will achieve 800 Gbps (Corporation, 2025e), rendering 6.4 Tbps bandwidth. As a reference, Google’s software transport Snap (Marty et al., 2019) can handle 80 Gbps traffic on a CPU core (though they do not use RDMA NICs). Our goal is to use 1 CPU core to handle 400G unidirectional traffic (i.e., 2 cores for 400G bidirectional traffic; excluding possible pacer cores for receiver-driven CC). To this end, we leverage the following techniques: